At least 99% of the data in any data set lie within how many standard deviation of the mean? Use Chebychov’s Theorem.
Answer #1

If we create an interval k= 10 standard deviations on either side of the mean chebyshavs theorem tells us

1 - (1 / k2)

= 1 - ( 1 / 102 )

= 1- (1/100 )

= 99/100

= 0.99

= 99%

Therefore

At least 99% of the data in any data set lie within 10 standard deviation of the mean